Numbers 1 to 10 are written on ten separate slips (one number on one slip), kept in a box and mixed well. One slip is chosen from the box without looking into it. What is the probability of . 
 getting a number greater than 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Number greater than 6 out of ten that are 7, 8, 9, 10. So there are 4possible outcomes.
∴, probability of getting a number greater than6 = 4/10 = 2/5

Thanks